Circular No. 8452

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

TRANSIENT IN LYNX E. J. Christensen, Lunar and Planetary Laboratory, reports the dis-

covery of a transient starlike object located at α = 8^{h}03^{m}24^{s}.60 ± 0^{s}.07, δ = +38^{o}18′35′′.9 ± 0′′.5 (equinox 2000.0). Nothing is visible within 0′.5 of this position on the Palomar Sky Survey or in the USNO-A2.0 catalogue. Following are red magnitudes derived from the CCD images, taken in the course of the Catalina Sky Survey with the Catalina 0.68-m Schmidt telescope: Dec. 11.4321 UT, [19.5; 11.4436, 15.9; 11.4549, 17.5; 11.4656, 18.2; 11.5073, 19.6; 11.5081, 18.9; 11.5090, 19.1; 11.5099, 19.7.

SUPERNOVAE 2004gq AND 2004gr Discoveries of two supernovae have been reported from unfiltered CCD

images: SN 2004gq independently by H. Pugh and W. Li (LOSS/KAIT; cf. IAUC 8448) and by F. Manzini (Novara, Italy, 0.4-m telescope; Stazione Astronomica di Sozzago Supernova Search), and SN 2004gr by LOSS.

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004gq Dec. 11.36 5^{h}12^{m}04.^{s}81 −15^{o}40′54′′.2 15.5 22′′.3 E, 22′′.4 N
2004gr Dec. 11.57 11 26 15.02 +27 52 06.7 17.9 9′′.2 W, 5′′.5 N

The data for SN 2004gq above are from LOSS; position end figures by R. Behrend (Geneva Observatory) from Manzini’s image: 04^{s}.87, 53′′.6. Manzini adds that nothing is visible at this location on Digitized Sky Survey images or on his earlier images (no dates specified). Additional magnitudes for SN 2004gq: Dec. 4.38 UT, [19.5 (KAIT); 11.93, 15.9 (Manzini); 12.02, 15.9 (Manzini); 12.30, 15.4 (KAIT). KAIT magnitudes for SN 2004gr: May 19.19, [19.5; Nov. 20.58, hint (poor conditions); Dec. 12.58, 17.9.

A. V. Filippenko and R. J. Foley, University of California, Berkeley, report that inspection of CCD spectra (range 320–940 nm), obtained on Dec. 12 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2004gq is probably of type Ic, roughly 4 days before maximum brightness. Its spectrum generally resembles the pre-maximum spectrum of SN 1994I (Filippenko et al. 1995, Ap.J. 450, L11), but there is an additional, weak absorption line at rest wavelength 635 nm that could be attributed Hα. The equivalent width of the narrow interstellar Na I D absorption line at 589 nm is ∼ 0.1 nm. Spectra of SN 2004gr show that it is of type II, probably the II-P variety and roughly a month past explosion, given the well-developed P-Cyg profiles of hydrogen Balmer, Fe II, and other lines.

2004 December 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green